# Cohet: A CXL-Driven Coherent Heterogeneous Computing Framework with Hardware-Calibrated Full-System Simulation

### v1

@@ VII Related Work
CXL simulation and emulation methods. Previous studies often used remote NUMA node (homogeneous in essence) for the emulation of CXL type-1/2 devices [huang_hal_2024, tarot_row_hammer, schuh_cc-nic_2024] or type-3 devices [wahlgren2022evaluating, hpc_cxl_memory, dmrpc, maruf2023tpp]. Although this approach has greatly facilitated early investigations, the performance characteristics of a remote NUMA node differ substantially from those of a genuine CXL device [demystifying_type2, cxl_memory_character], which can in some cases lead to misleading conclusions [demystifying_type3]. Existing software simulators focus solely on modeling memory expansion, pooling, and sharing for type-3 devices [QEMU, mess, CXLMemSim_github, cxl_dmsim], without any touch on heterogeneous systems with typ-1/2 devices. Table II compares SimCXL with other existing CXL system simulation and emulation tools across several dimensions. Only SimCXL models CXL XPU devices and all three CXL sub-protocols, and supports full-system simulation of Cohet computing framework. After meticulous hardware calibration, SimCXL achieves a low simulation error of 3%.
CXL memory disaggregation and pooling. A wealth of prior work has evaluated real CXL type-3 devices, providing comprehensive performance characterizations and practical insights [demystifying_type3, exploring_cxl_asic, cxl_memory_character, samsung_cxl_ssd]. Many studies have also explored CXL-based tiered memory management at the OS level. [maruf2023tpp, m5, tiered_memory, nomad, neomem]. In addition, some studies exploit CXL memory to expand system memory capacity and bandwidth to accelerate memory-intensive applications [ahnexamination, recomm, hpc_cxl_memory, gpu_graph_process, lia_cxl_llm].
CXL-based heterogeneous computing. Recently, a study provides a detailed performance characterization of the CXL type-2 device (FPGA) and offloads two OS kernel memory optimization functions [demystifying_type2]. Some studies leverage the cache coherence and unified memory view provided by CXL.cache to optimize specific application scenarios [huang_hal_2024, tarot_row_hammer, xu_efficient_tensor]. Other works explore host-accelerator synchronization mechanisms based on cache coherence [schuh_cc-nic_2024, mozart, rambda, rag_cxl]. However, these works are limited to: 1) hardware customization exploiting selected CXL.cache features with direct manipulations on physical memory addresses (FPGA-like development), 2) predominately relying on remote NUMA emulation for evaluation. In contrast, our work builds a Cohet computing framework with a global perspective on the software stack and hardware architecture, greatly simplifying programming model and memory management. We also introduce SimCXL, a research tool that fills a critical gap in CXL-enable heterogeneous systems simulation infrastructure.
@@ VIII Discussion and Future Work
Despite its promise, CXL-driven coherent heterogeneous computing still faces many challenges. The overhead of ATS-based address translation remains unexplored due to the lack of ATS support on current CXL FPGA platforms [no_support_ats]; prior CCIX studies indicate substantial ATC miss penalties [evaluation_ccix]. Furthermore, OS support for CXL devices remains immature [cxl_linux], and mainstream programming frameworks like OpenCL have not yet integrated CXL workflows.
Our future work will focus on adding new features to SimCXL, currently limited to CXL 1.1. With an industry target toward CXL 3.x and its fabric-based multi-node interconnects [CXL_spec], we are designing CXL switch models to extend SimCXL and exploring integration with SimBricks [simbricks] for large-scale fabric-based system simulation. As the coherence domain scales up (i.e., with more child nodes in a supernode), the performance loss and design complexity of maintaining hardware coherence increase sharply, and the huge coherence traffic can become a system bottleneck. To mitigate coherence-traffic storms, we plan to explore a hierarchical coherence protocol for small-scale supernodes. Each child node interacts with a local agent for coherence transactions; the local agent consults a global agent only if it lacks the requested replica.
Furthermore, we will continue to explore the broader applications of Cohet. In graph processing, large datasets can be stored in a unified memory pool, and graph algorithms with fine-grained random-access patterns [graph1, graph2] offloaded to CXL accelerators can benefit from the coherent CXL interconnect. Similarly, in-memory key-value store operations (e.g., GET/PUT) offloaded to CXL accelerators will benefit from lower-latency, fine-grained memory accesses. Moreover, the superior attributes of CXL such as memory pooling and sharing, p2p direct memory access, and TEE security offer compelling opportunities to accelerate workloads such as large language model inference and distributed services.

### v2

@@ VII Related Work
CXL simulation and emulation methods. Previous studies often used remote NUMA node (homogeneous in essence) for the emulation of CXL type-1/2 devices [33, 74, 67] or type-3 devices [25, 84, 97, 51]. Although this approach has greatly facilitated early investigations, the performance characteristics of a remote NUMA node differ substantially from those of a genuine CXL device [36, 47], which can in some cases lead to misleading conclusions [78]. Existing software simulators focus solely on modeling memory expansion, pooling, and sharing for type-3 devices [20, 21, 92, 86], without any touch on heterogeneous systems with typ-1/2 devices. Table II compares SimCXL with other existing CXL system simulation and emulation tools across several dimensions. Only SimCXL models CXL XPU devices and all three CXL sub-protocols, and supports full-system simulation of Cohet computing framework. After meticulous hardware calibration, SimCXL achieves a low simulation error of 3%.
CXL memory disaggregation and pooling. A wealth of prior work has evaluated real CXL type-3 devices, providing comprehensive performance characterizations and practical insights [78, 81, 47, 96]. Many studies have also explored CXL-based tiered memory management at the OS level. [51, 77, 83, 90, 100]. In addition, some studies exploit CXL memory to expand system memory capacity and bandwidth to accelerate memory-intensive applications [3, 46, 84, 65, 41].
CXL-based heterogeneous computing. Recently, a study provides a detailed performance characterization of the CXL type-2 device (FPGA) and offloads two OS kernel memory optimization functions [36]. Some studies leverage the cache coherence and unified memory view provided by CXL.cache to optimize specific application scenarios [33, 74, 91]. Other works explore host-accelerator synchronization mechanisms based on cache coherence [67, 79, 94, 64]. However, these works are limited to: 1) hardware customization exploiting selected CXL.cache features with direct manipulations on physical memory addresses (FPGA-like development), 2) predominately relying on remote NUMA emulation for evaluation. In contrast, our work builds a Cohet computing framework with a global perspective on the software stack and hardware architecture, greatly simplifying programming model and memory management. We also introduce SimCXL, a research tool that fills a critical gap in CXL-enable heterogeneous systems simulation infrastructure.
@@ VIII Discussion and Future Work
Despite its promise, CXL-driven coherent heterogeneous computing still faces many challenges. The overhead of ATS-based address translation remains unexplored due to the lack of ATS support on current CXL FPGA platforms [87]; prior CCIX studies indicate substantial ATC miss penalties [80]. Furthermore, OS support for CXL devices remains immature [16], and mainstream programming frameworks like OpenCL have not yet integrated CXL workflows.
Our future work will focus on adding new features to SimCXL, currently limited to CXL 1.1. With an industry target toward CXL 3.x and its fabric-based multi-node interconnects [18], we are designing CXL switch models to extend SimCXL and exploring integration with SimBricks [43] for large-scale fabric-based system simulation. As the coherence domain scales up (i.e., with more child nodes in a supernode), the performance loss and design complexity of maintaining hardware coherence increase sharply, and the huge coherence traffic can become a system bottleneck. To mitigate coherence-traffic storms, we plan to explore a hierarchical coherence protocol for small-scale supernodes. Each child node interacts with a local agent for coherence transactions; the local agent consults a global agent only if it lacks the requested replica.
Furthermore, we will continue to explore the broader applications of Cohet. In graph processing, large datasets can be stored in a unified memory pool, and graph algorithms with fine-grained random-access patterns [71, 93] offloaded to CXL accelerators can benefit from the coherent CXL interconnect. Similarly, in-memory key-value store operations (e.g., GET/PUT) offloaded to CXL accelerators will benefit from lower-latency, fine-grained memory accesses. Moreover, the superior attributes of CXL such as memory pooling and sharing, p2p direct memory access, and TEE security offer compelling opportunities to accelerate workloads such as large language model inference and distributed services.